OpenStudy (anonymous):

show work and simplify -4√(x+9)=20

hero (hero):

√(x+9)= -5 x + 9 = 25 x = 16

OpenStudy (anonymous):

thats wrong , i got no solution

hero (hero):

Yeah, that's probably right

OpenStudy (anonymous):

you forgot to use the -4

hero (hero):

I did use the negative 4

hero (hero):

How do you think I got -5?

OpenStudy (anonymous):

thats -5

hero (hero):

-4√(x+9)= 20 √(x+9)= -5 x + 9 = 25 x = 16

hero (hero):

Those were my original steps

OpenStudy (anonymous):

well it might be right !

hero (hero):

I think it is correct. I solved for x

hero (hero):

Nah, I checked on wolfram alpha. It's no solutions